OpenStudy (anonymous):

solve the logarithmic equation. ln (x-8) - ln (x+7) = ln (x-10) - ln (x+8)

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

use the property. ln(a)-ln(b)= ln(a/b)

OpenStudy (anonymous):

how did you get to -2? thats where i got stuck at

OpenStudy (anonymous):

after u use the property, u note that: (x-8)/(x+7)=(x-10)/(x+8). solve this equation to get x

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no prob.